OpenStudy (anonymous):

An airplane starting from airport A flies 300 km east, then 340 km at 32.0° west of north, and then 130 km north to arrive finally at airport B. How far will the pilot travel in this direct flight? At what angle?

OpenStudy (anonymous):

Draw a picture of it, including the north and west components of the one with the angle. Then you can add or subtract all the east, west components, to get the 'x' component. And add up the north components, to get the 'y' component. Then the total distance is: (x^2 + y^2)^1/2 To get the angle, use: tan(theta) = y/x

OpenStudy (anonymous):

Here are the two equations I wrote above: \[d = \sqrt{x^2 + y^2}\] \[\tan\theta = \frac{y}{x}\]